the perimeter of a rectangle is 86 feet. the width is 6 feet less than the length.what is the length and width of the rectangle?

Answer by ewatrrr(24785)

 
Hi,
.
L + w = 43
L + (L-6) = 43
2L = 49
L = 24.5ft and w = 18.5ft 24.5+-+6
and...checking
24.5 + 18.5 = = 43